OpenStudy (anonymous):

HELP!! Simplify each and state the excluded values. M-8/8m-64, n-2/10n-20, n^2+18+81/n^2+17n+72, 12x+24/30x+6

OpenStudy (anonymous):

***n^2+18n+81

OpenStudy (freckles):

are there any type-o's?

OpenStudy (anonymous):

Yeah on the third one

OpenStudy (freckles):

\[m-\frac{8}{8}m-64 ? \\ n-\frac{2}{10}n-20? \\ n^2+18n+\frac{81}{n^2}+17n+72 ? \\ 12x+\frac{24}{30}x+6?\]

OpenStudy (freckles):

this looks funny to me

OpenStudy (freckles):

I think you meant (m-8)/(8m-64) (n-2)/(10n-20) (n^2+18n+81)/(n^2+17n+72) (12x+24)/(30x+6)

OpenStudy (freckles):

which has translation \[\frac{m-8}{8m-64} \\ \frac{n-2}{10n-20} \\ \frac{n^2+18n+81}{n^2+17n+72} \\ \frac{12x+24}{30x+6}\]

OpenStudy (freckles):

anyways to find the excluded values set all the bottoms of any fractions equal to zero and solve